# Molecular Fingerprints Are Strong Models for Peptide Function Prediction

### v1

@@ 5 Conclusions
We presented an approach to peptide function prediction using count-based variants of hashed molecular fingerprints. They utilize atomic-level representation of peptides using their molecular graphs, which is cheap to obtain. This feature extraction method is also extremely fast to compute. Despite being strictly local, short-range, and generally simple descriptors, they show remarkable performance across a wide range of tasks.
Combined with the LightGBM classifier, this approach achieves superior results in a comprehensive evaluation of six benchmarks consisting of 126 datasets. In addition to classification quality, they are highly performant and parameter-efficient. They outperform a wide variety of other approaches, including GNNs, pretrained amino acid sequence transformers, feature engineering, and multimodal models.
This work also constitutes, to the best of our knowledge, the largest comparison of fingerprints applied to peptides in the literature. In contrast to previous bioinformatics works, we use count variants, instead of binary fingerprints, and show that this detail is crucial to obtain a strong encoding. Our major methodological contribution is showing that these inherently short-range models obtain SOTA results on peptide tasks from Long Range Graph Benchmark (LRGB). ECFP-based classifier with radius 2 outperforms the previous best model by 1.5% AUPRC on Peptides-func. These results challenge the assumed importance of long-range dependencies on this task, further strengthening previous observations from the literature (Tönshoff et al., 2024).
@@
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional benchmark metrics
Here, we report additional metrics for AMPs benchmarks from (Gao et al., 2024) (Table 8) and (Xu et al., 2021) (Table 9), as well as AutoPeptideML benchmark (Fernández-Díaz et al., 2024) (Table 10). They were omitted in the main text due to space limitations.
Table 8: Results for BERT benchmark (Gao et al., 2024) results.
Metric Model ADAPTABLE APD CAMP dbAMP DRAMP YADAMP avg
Recall AMP-BERT 67.60 79.40 80.40 54.20 54.70 82.60 69.82
Bert-Protein 58.40 94.10 76.30 83.80 47.60 47.60 67.97
cAMPs_pred 44.50 58.80 57.90 36.30 24.70 73.90 49.35
LM_pred 40.00 55.90 49.10 45.10 31.80 69.60 48.58
LM_pred (BFD) 64.60 88.20 78.00 58.50 52.80 82.60 59.67
ECFP 75.60 97.10 88.70 95.10 66.80 100.00 87.22
$T$ 75.60 97.10 89.10 94.00 67.10 100.00 87.15
RDKit 77.20 97.10 88.70 94.00 66.50 100.00 87.25
Precision AMP-BERT 80.20 77.10 84.20 74.40 77.10 82.60 79.27
Bert-Protein 79.20 86.50 85.40 82.10 77.10 77.10 81.23
cAMPs_pred 98.20 100.00 97.90 94.50 94.80 100.00 97.57
LM_pred 92.50 100.00 94.00 92.80 94.00 100.00 95.55
LM_pred (BFD) 86.30 91.20 90.20 83.40 83.90 86.40 91.62
ECFP 86.90 89.20 87.70 88.20 89.00 92.00 88.83
$T$ 86.40 91.70 88.50 87.80 89.80 92.00 89.37
RDKit 87.50 89.20 88.00 89.00 89.80 85.20 88.12
AUROC AMP-BERT 81.70 91.80 90.60 78.40 76.90 90.70 85.02
Bert-Protein 79.40 96.10 90.90 91.90 74.00 74.00 84.38
cAMPs_pred 82.30 91.80 87.00 81.90 68.10 93.00 84.02
LM_pred 88.00 97.20 92.30 89.50 82.80 98.70 91.42
LM_pred (BFD) 86.30 97.40 93.60 87.60 83.90 93.40 90.47
ECFP 89.50 98.40 94.30 97.20 87.80 100.00 94.53
$T$ 89.30 98.20 94.30 97.10 87.60 100.00 94.42
RDKit 90.20 99.30 94.80 97.00 88.20 100.00 94.92
F1 AMP-BERT 81.70 78.30 82.30 62.70 64.00 82.60 75.27
Bert-Protein 79.40 90.10 80.60 82.90 58.80 58.80 75.10
cAMPs_pred 61.30 74.10 72.80 52.40 39.20 85.00 64.13
LM_pred 55.90 71.70 64.50 60.70 47.60 82.10 63.75
LM_pred (BFD) 73.90 89.60 83.60 68.70 64.80 84.40 70.32
ECFP 80.80 93.00 88.20 91.50 76.30 95.80 87.60
$T$ 80.60 94.30 88.80 90.80 76.90 95.80 87.87
RDKit 82.00 93.00 88.40 91.40 76.40 92.00 87.20
Table 9: Additional metrics for results on (Xu et al., 2021).
Method Accuracy AUROC F1 MCC Recall Specificity
AMPscannerV2 56.8 58.5 54.8 0.137 52.3 61.3
iAMP-2L 59.2 59.2 36.8 0.261 23.8 94.7
ADAM-SVM 61.2 61.2 47.1 0.264 34.6 87.8
ampir 56.3 61.9 37.9 0.156 26.6 85.9
MLAMP 55.3 62.9 23 0.194 13.3 97.2
ADAM-HMM 68.4 68.4 62.3 0.39 52.1 84.7
AMPlify 64.2 69.7 46.2 0.381 30.7 97.6
AMPEP 65.8 72.7 48.7 0.425 32.5 99.2
AMPfun 67.4 73.5 55.5 0.414 40.6 94.3
RDKit 67.5 73.7 58.5 0.388 45.8 89.1
ECFP 69.3 75.3 60.9 0.426 47.9 90.6
$T$ 69.4 74.9 61.7 0.424 49.3 89.5
Table 10: Additional metrics for results on AutoPeptideML (Fernández-Díaz et al., 2024).
Model # params Accuracy MCC AUROC F1
ProtBERT 420M 68.5 0.375 75.9 67.7
ESM2-150M 150M 69.8 0.402 77.7 68.5
Prost-T5 3B 70.0 0.409 77.1 69.0
ESM2-8M 8M 70.2 0.418 77.5 69.4
ESM2-35M 35M 71.0 0.428 78.0 70.2
ESM1b-650M 650M 71.1 0.433 78.9 69.2
ESM2-650M 650M 68.0 0.438 79.7 69.7
Prot-T5-XL 3B 68.9 0.447 79.7 70.4
RDKit 20k 70.7 0.421 76.9 69.6
$T$ 23k 70.8 0.422 77.1 69.4
ECFP 22k 71.5 0.437 78.1 70.2
Appendix B Additional classifiers on LRGB
Here, we present results of two additional classifiers on the LRGB benchmark datasets: Random Forest (Breiman, 2001) and Extremely Randomized Trees (Geurts et al., 2006). For both we use 500 trees, entropy or squared error as cost function (for classification and regression, respectively), and class weighting for classification. We can report standard deviations for them over 10 random seeds, because they are nondeterministic.
Peptides-func results are in Table 11, and Peptides-struct in Table 12. For clarity, we also include LightGBM results from the main body, which is a deterministic classifier and thus does not have standard deviation.
Random Forest and Extremely Randomized Trees classifiers have very low standard deviation, highlighting the stability of the fingerprint-based approach. Models based on ECFP features consistently achieve the best results in all cases, but other fingerprints and classifiers also obtain results highly competitive with long-range GNNs.
Table 11: AUPRC $\uparrow$ for different classifiers on Peptides-func.
LightGBM Random Forest Extremely Randomized Trees
[…]
Table 12: MAE $\downarrow$ for different classifiers on Peptides-struct.
LightGBM Random Forest Extremely Randomized Trees
[…]
Appendix C Additional LRGB timings
Here, we present expanded timings of fingerprints and compare them to times reported in LRGB (Dwivedi et al., 2022) (Appendix C.2), summarizing them in Table 13. Times for fingerprint-based models are an average of 10 runs on 12-core Intel Core i7-12700KF. We measure them for Peptides-func, and for Peptides-struct results were almost identical. Timings for GNNs are taken from (Dwivedi et al., 2022), and use Nvidia A100 GPU.
The entire time for our approach (feature extraction + classifier training) is shorter than precomputing structural embeddings or training even a single epoch of SAN model. Graph transformer model has faster epochs, they still require precomputing LapPE embeddings, while also giving much worse results than fingerprints (see main body for results table).
One should also take into consideration the difference in raw compute power (CPU vs powerful GPU) required to get those times, which makes the difference even more significant.
Table 13: Time of computation on Peptides-func.
Model Time [s]
Transformer+LapPE 5.9
SAN+LapPE 53.6
SAN+RWSE 49.7
LapPE encoding 74
RWSE encoding 53
ECFP 19
$T$ 15.2
RDKit 42.8
Appendix D Additional PeptideReactor results
Here, we present the performance of all encodings on PeptideReactor (Spänig et al., 2021), in Table 14.
Table 14: Full results on PeptideReactor benchmark (Spänig et al., 2021).
Encoding Avg F1 Type
FP encoding 82.9 fingerprint
cksaap 82.4 sequence
dist_f 82.2 sequence
psekraac 81.8 sequence
ECFP 81.7 fingerprint
ngram_ 81.6 sequence
dde 81.3 sequence
dpc 80.6 sequence
$T$ 80.6 fingerprint
RDKit 80.6 fingerprint
fldpc_ 79.4 sequence
qsorde 78.9 sequence
waac_a 78.8 sequence
apaac_ 78.2 sequence
aac 78.1 sequence
binary 77.9 sequence
ctdd 77.9 sequence
paac_l 77.9 sequence
aainde 77.8 sequence
tpc 76.6 sequence
cksaag 76.5 sequence
ctriad 75.9 sequence
ctdt 75.7 sequence
ctdc 75.6 sequence
ksctri 75.6 sequence
gtpc 75 sequence
gdpc 74 sequence
flgc_a 73.8 sequence
fft_aa 73.3 sequence
qsar 72.8 structure
nmbrot 72.3 sequence
zscale 72.1 sequence
eaac_w 71.6 sequence
blomap 70.9 sequence
egaac_ 70.3 sequence
blosum 69.8 sequence
delaun 69.5 structure
gaac 69.5 sequence
moran_ 69 sequence
socnum 69 sequence
geary_ 68.9 sequence
cgr_re 68.8 sequence
distan 61.3 structure
electr 60.9 structure
disord 53.5 structure
sseb 51.7 structure
asa 51.5 structure
ta 51.4 structure
ssec 48.3 structure

### v2

@@ 7 Conclusions
We presented an approach to peptide function prediction using count-based variants of hashed molecular fingerprints. They utilize atomic-level representation of peptides using their molecular graphs, which are inexpensive to obtain. This feature extraction method is also extremely fast to compute and does not require knowledge of the folded peptide structure. Despite being strictly local, short-range, and generally simple descriptors, they show remarkable performance across a wide range of tasks.
Combined with the LightGBM classifier, this approach achieves superior results in a comprehensive evaluation of six benchmarks consisting of 132 datasets. In addition to classification quality, they are highly performant and parameter-efficient. They outperform a wide variety of other approaches, including GNNs, pretrained amino acid sequence transformers, feature engineering, and multimodal models.
This work also constitutes, to the best of our knowledge, the largest comparison of fingerprints applied to peptides in the literature. In contrast to previous bioinformatics works, we use count variants, instead of binary fingerprints, and show that this detail is crucial to obtain strong encodings. Our major methodological contribution is showing that these inherently short-range models obtain SOTA results on peptide tasks from Long Range Graph Benchmark (LRGB). ECFP-based classifier with radius 2 outperforms the previous best model by 1.5% AUPRC on Peptides-func. These results challenge the assumed importance of long-range dependencies on this task, further strengthening previous observations from the literature LRGB_reassessment .
@@
8 Acknowledgements
We thank BIT Student Scientific Group for computational resources and aid for this project. We thank Aleksandra Elbakyan for her work and support for accessibility of science.
8.1 Data availability
All datasets and benchmarks used are publicly available at https://github.com/scikit-fingerprints/peptides_molecular_fingerprints_classification.
9 Author contributions statement
Contributions follow the CRediT format. J.A. was responsible for conceptualization, methodology, software, validation, investigation, data curation, and writing. P.L. was responsible for software, investigation, and writing. W.C. was responsible for writing, review, and supervision.
10 Competing interests
No competing interest is declared.
11 Funding
This work was supported by the general funding from the Ministry of Science and Higher Education to AGH University in Krakow, and by project Excellence Initiative – Research University (IDUB) for AGH University of Krakow.